Which of the following mass measurements is the smallest?
a. 78 kg
b. 78 mg
c. 78 pg
d. 78 fg
e. 78 ng
Answer #1

(A)1 Kg=1000g

78 Kg=78*1000=78000g

(B)1 mg=0.001g

78mg=78*0.001g=0.078g

(C)1pg=1e-12g

78pg=78*(1e-12)=(7.8e-11)g=10.202g

(D)1fg=(1e-15)g

78fg=78*(1e-15)=7.800000000000001e-14=7.202g

(E)1ng=1e-9g

78ng=78*(1e-9)=7.8e-8=21.202-9=12.202g

so,(B) has less mass measurement.Hence (B) is the correct answer